# Denoising Complex Covariance Matrices with Hybrid ResNet and Random Matrix Theory: Cryptocurrency Portfolio Applications

### v1

@@ 4 Covariance models
    where $\mathbf{L}$ is a rectangular matrix of dimension $p\times k$, with $k$ denoting the number of blocks. In particular, we set $p=100$ and $k=12$, with heterogeneous block sizes given by $[3,3,4,5,6,7,7,9,11,13,15,17]$, but with equal intra-block correlation intensity $\gamma=0.3$. To ensure a proper correlation matrix, the diagonal entries are set to one. Figure 2(a) displays the population model under these specifications (left) and a finite-sample realization with $n=200$ (right). This block-independent and homogeneous structure model has been previously examined in [16].
@@
    Here, we set $\gamma=0.1$. Figure 2(b) presents the population model under these specifications (left) with $p=100$, and a finite-sample realization with $n=200$ (right). This system was originally proposed in [17] as a model to characterize the complex interaction structure of financial markets. In particular, the population eigenvalues of this system correspond to the solutions of a symmetric tridiagonal matrix and exhibit deep mathematical connections with Fibonacci and Lucas numbers [30].
@@
Figure 3 presents the log-log scree plot of the eigenvalues for each model. Model 1 exhibits $k=12$ non-degenerate eigenvalues, whose magnitudes correspond to the size of each of the $k$ blocks (see [16]). Model 2 displays the intriguing property that its eigenvalue spectrum approximately follows a power-law decay. Due to this property, it has been conjectured in [17] that it can capture the stylized facts underlying the complex interactions of financial markets. In the case of Model 3, the power-law behavior is imposed by construction with a slope of $\alpha=1.5$, providing a further step toward characterizing the complex interactions of financial markets through their covariance structure.
@@ 6 Empirical data
In Figure 5(a), we present the empirical covariance matrix of the cryptocurrency dataset. The assets have been ordered using the seriation algorithm proposed by [31] with a twofold objective: (i) to enhance the graphical representation by positioning assets with similar covariance patterns closer to each other, and (ii) to improve the ability of deep learning estimators to detect and learn the structural dependencies within the noisy covariance matrices. The resulting ordered covariance matrix exhibits patterns that closely resemble those generated by the power-law model, suggesting that the empirical structure of the cryptocurrency market may share similar scaling properties. Figure 5(b) displays the scree plot of the eigenvalues of the empirical covariance matrix. The observed behavior closely resembles that of Models 2 and 3, with the exception of a few outliers in the tails. Moreover, the dendrogram presented in Figure 5(c) exhibits a hierarchical pattern qualitatively consistent with the structures observed in these two models. These results indicate that our proposed covariance models effectively capture the stylized facts present in the empirical cryptocurrency dataset.
Figure 5: Cryptocurrency empirical data (a) Covariance matrix of ordered elements. (b) Scree plot of eigenvalues. (c) Dendrogram under the same methodology as the hierarchical estimator with single linkage (SLCA). For illustrative purpose the descendant links below a cluster node $k$ are equally colored if $k$ is the first node below the cut threshold $t(=0.7)$
@@
Using a stride of 1, we thus obtain a training sample of 100 datasets, each of dimension $p\times N$, to apply the learning process. The ResNet algorithm is configured using the same settings as in the simulation study. Next, portfolio returns are computed by multiplying the in-sample allocations $\mathbf{w}_{in}$ with the out-of-sample returns $\mathbf{R}_{out}$. Figure 6(a) displays the cumulative returns for each estimator throughout the walk-forward analysis, with the rebalancing dates indicated by grey dashed vertical lines. Figure 6(b) shows the performance of the top 12 individual cryptocurrencies over the same period. The returns of individual cryptocurrencies exhibit greater volatility, reflecting higher risk, and consequently they can achieve higher cumulative returns compared to the diversified portfolios.
@@
Estimator Cumulative Return Annual Return Annual Volatility Sharpe Ratio Maximum Drawdown Turnover
@@
The annual portfolio volatility is measured by eq. 6 and also multiplied by the factor 365 to obtain a proxy. Interestingly, the naive estimator reaches the minimum value under this metric. Thus, an increase in mean return does not necessarily imply a decrease in standard deviation. The Sharpe Ratio, defined as the ratio between the mean and standard deviation of the portfolio return, gives us a fairer number to compare different investments. Under this measure, the most equilibrated portfolio is obtained again by the $\mathbf{\Xi}^{2S(CNN)}$ estimator.
@@
Figure 5 shows the equivalent financial metrics but for the buy & hold strategy of investing in individual cryptocurrencies through the walk-forward period. Notice that here, the turnover metric is not measured because we never rebalance. Interestingly, several coins surpass the performance of the combination of a state-of-the-art covariance estimator and the classical investment strategy MVP+. Even accounting for the trade-off measured by the Sharpe Ratio and the Maximum Drawdown, crypto like GT-USD outperforms all strategies. Therefore, an important direction for future research is to integrate asset selection into the overall covariance estimation and portfolio allocation framework.
Cryptocurrency Cumulative Return Annual Return Annual Volatility Sharpe Ratio Maximum Drawdown
@@ 7 Conclusion
We found that the two-step estimator based on the ResNet architecture ($\mathbf{\Xi}^{2S(CNN)}$) minimizes the Frobenius loss function for the block-diagonal model. For the same model, the two-step hybrid estimator ($\mathbf{\Xi}^{2S(H)}$) attains the minimum value of the MV loss function. As the complexity of the covariance model increases, we observe that applying the second filtering step becomes unnecessary; in such cases, the single-step estimators ($\mathbf{\Xi}^{CNN}$) and ($\mathbf{\Xi}^{H}$) are sufficient to minimize noise with respect to the Frobenius and MV losses, respectively. On the other hand, an interesting phenomenon arises when the covariance model follows a power-law structure: in this case, none of the state-of-the-art estimators outperform the naive one in terms of the Frobenius loss. However, regarding the MV loss, the two-step hybrid estimator achieves the lowest noise level. Overall, a consistent pattern emerges MV loss systematically decreases across all models when using the ResNet-based estimators. The distinctive point here is that the MV loss captures deviations along the principal directions of variance and, as such, is more suitable for portfolio optimization applications.
@@
On the other hand, when considering empirical data from the cryptocurrency market, we observe that the covariance structure, the scree plot, and the dendrogram exhibit stylized facts similar to those found in both the fully hierarchical model and the power-law model. Thus, we initially confirm the conjecture proposed in [17], which states that the first model adequately characterizes the complex dynamics of financial instruments—this time applied to a set of cryptocurrencies. Furthermore, we extend this conjecture to a new model based on power-law behavior.
Focusing on the specific dataset analyzed, we find that the best financial metrics were generally obtained using the $(\mathbf{\Xi}^{2S(CNN)})$ estimator within the MVP+ portfolio framework. This result holds even under the challenging configuration of our experiment, where the training set corresponded to a bull market period, while the testing set covered a heterogeneous phase that began with a bear market trend.
@@
Nevertheless, although these results are promising in terms of adequately characterizing the complex dynamics of cryptocurrencies and proposing a hybrid estimator that combines the advantages of state-of-the-art neural network models with the analytical foundations of random matrix theory, they are not sufficient to achieve the maximum cumulative return. When comparing with the performance of individual assets, we find that there are instruments capable of nearly tripling the initial capital while offering a better balance between risk exposure and profitability (Sharpe Ratio). Therefore, from a practical standpoint, there is still room for improvement in the search for more profitable strategies. In this regard, it would be interesting to continue combining methodologies and to incorporate the last missing ingredient—one that is uncommon within the econophysics field: asset selection and valuation from a fundamental analysis perspective [32]. However, the challenge is substantial, as cryptocurrencies generally lack traditional fundamental metrics available in conventional financial markets.
Declaration of generative AI and AI-assisted technologies in the manuscript preparation process.
@@
During the preparation of this work, the author used ChatGPT to assist with readability improvements and proofreading. After using this tool/service, the author reviewed and edited the content as needed and takes full responsibility for the content of the published article.

### v2

@@ 4 Covariance models
    where $\mathbf{L}$ is a rectangular matrix of dimension $p\times k$, with $k$ denoting the number of blocks. In particular, we set $p=100$ and $k=12$, with heterogeneous block sizes given by $[3,3,4,5,6,7,7,9,11,13,15,17]$, but with equal intra-block correlation intensity $\gamma=0.3$. To ensure a proper correlation matrix, the diagonal entries are set to one. Figure 2(a) displays the population model under these specifications (left) and a finite-sample realization with $n=200$ (right). This block-independent and homogeneous structure model has been previously examined in [garcia2023two].
@@
    Here, we set $\gamma=0.1$. Figure 2(b) presents the population model under these specifications (left) with $p=100$, and a finite-sample realization with $n=200$ (right). This system was originally proposed in [garcia2025high] as a model to characterize the complex interaction structure of financial markets. In particular, the population eigenvalues of this system correspond to the solutions of a symmetric tridiagonal matrix and exhibit deep mathematical connections with Fibonacci and Lucas numbers [cahill2004fibonacci].
@@
Figure 3 presents the log-log scree plot of the eigenvalues for each model. Model 1 exhibits $k=12$ non-degenerate eigenvalues, whose magnitudes correspond to the size of each of the $k$ blocks (see [garcia2023two]). Model 2 displays the intriguing property that its eigenvalue spectrum approximately follows a power-law decay. Due to this property, it has been conjectured in [garcia2025high] that it can capture the stylized facts underlying the complex interactions of financial markets. In the case of Model 3, the power-law behavior is imposed by construction with a slope of $\alpha=1.5$, providing a further step toward characterizing the complex interactions of financial markets through their covariance structure.
@@ 6 Empirical data
In Figure 5(a), we present the empirical covariance matrix of the cryptocurrency dataset. The assets have been ordered using the seriation algorithm proposed by [atkins1998spectral] with a twofold objective: (i) to enhance the graphical representation by positioning assets with similar covariance patterns closer to each other, and (ii) to improve the ability of deep learning estimators to detect and learn the structural dependencies within the noisy covariance matrices. The resulting ordered covariance matrix exhibits patterns that closely resemble those generated by the power-law model, suggesting that the empirical structure of the cryptocurrency market may share similar scaling properties. Figure 5(b) displays the scree plot of the eigenvalues of the empirical covariance matrix. The observed behavior closely resembles that of Models 2 and 3, with the exception of a few outliers in the tails. The fitted slope of $\alpha=0.2$ is superimposed as a graphical reference. Moreover, the dendrogram presented in Figure 5(c) exhibits a hierarchical pattern qualitatively consistent with the structures observed in these two models. These results indicate that our proposed covariance models effectively capture the stylized facts present in the empirical cryptocurrency dataset.
Figure 5: Cryptocurrency empirical data (a) Covariance matrix of ordered elements. (b) Scree plot of eigenvalues with a fitted slope of $\alpha=0.2$. (c) Dendrogram under the same methodology as the hierarchical estimator with single linkage (SLCA). For illustrative purpose the descendant links below a cluster node $k$ are equally colored if $k$ is the first node below the cut threshold $t(=0.7)$
@@
Using a stride of 1, we thus obtain a training sample of 100 datasets, each of dimension $p\times N$, to apply the learning process. The ResNet algorithm is configured using the same settings as in the simulation study. Next, out-of-sample portfolio returns ( $\mathbf{R}_{\text{out}}$) are computed by multiplying the in-sample allocations ($\mathbf{w}_{\text{in}}^{\top}$) with the out-of-sample returns ($\mathbf{r}_{\text{out}}$):
$\mathbf{R}_{t}^{\text{out}}=\mathbf{w}_{\text{in}}^{\top}\mathbf{r}_{\text{out}}.$ (28)
Figure 6(a) displays the cumulative returns for each estimator throughout the walk-forward analysis, with the rebalancing dates indicated by grey dashed vertical lines. Figure 6(b) shows the performance of the top 12 individual cryptocurrencies over the same period. The returns of individual cryptocurrencies exhibit greater volatility, reflecting higher risk, and consequently they can achieve higher cumulative returns compared to the diversified portfolios.
@@
Estimator Cumulative Return Annual Return Annual Volatility SR Maximum Drawdown Turnover
@@
The annual portfolio volatility is measured by eq. 6 and also multiplied by the factor 365 to obtain a proxy. Interestingly, the naive estimator reaches the minimum value under this metric. Thus, an increase in mean return does not necessarily imply a decrease in standard deviation. The Sharpe Ratio (SR), defined as the ratio between the mean and standard deviation of the portfolio return, gives us a fairer number to compare different investments. Under this measure, the most equilibrated portfolio is obtained again by the $\mathbf{\Xi}^{2S(CNN)}$ estimator.
@@
Figure 5 shows the equivalent financial metrics but for the buy & hold strategy of investing in individual cryptocurrencies through the walk-forward period. Notice that here, the turnover metric is not measured because we never rebalance. Interestingly, several coins surpass the performance of the combination of a state-of-the-art covariance estimator and the classical investment strategy MVP+. Even accounting for the trade-off measured by the SR and the Maximum Drawdown, crypto like GT-USD outperforms all strategies. Therefore, an important direction for future research is to integrate asset selection into the overall covariance estimation and portfolio allocation framework.
Cryptocurrency Cumulative Return Annual Return Annual Volatility SR Maximum Drawdown
@@ 7 Conclusion
We found that the two-step estimator based on the ResNet architecture ($\mathbf{\Xi}^{2S(CNN)}$) minimizes the Frobenius loss function for the block-diagonal model. For the same model, the two-step hybrid estimator ($\mathbf{\Xi}^{2S(H)}$) attains the minimum value of the MV loss function. As the complexity of the covariance model increases, we observe that applying the second filtering step becomes unnecessary; in such cases, the single-step estimators ($\mathbf{\Xi}^{CNN}$) and ($\mathbf{\Xi}^{H}$) are sufficient to minimize noise with respect to the Frobenius and MV losses, respectively. Moreover, an interesting phenomenon arises when the covariance model follows a power-law structure: in this case, none of the state-of-the-art estimators outperform the naive one in terms of the Frobenius loss. However, regarding the MV loss, the two-step hybrid estimator achieves the lowest noise level. Overall, a consistent pattern emerges MV loss systematically decreases across all models when using the ResNet-based estimators. The distinctive point here is that the MV loss captures deviations along the principal directions of variance and, as such, is more suitable for portfolio optimization applications. In sum, as market structure becomes more complex (hierarchical or power-law), hybrid estimators increasingly outperform traditional methods in terms of portfolio risk control.
@@
On the other hand, when considering empirical data from the cryptocurrency market, we observe that the covariance structure, the associated scree plot, and the dendrogram, exhibit stylized facts closely resembling those found in both the nested hierarchical covariance model and the power-law covariance model. In particular, the eigenvalue spectrum displays an approximate power-law decay with coefficient $\alpha=0.2$. These observations provide initial confirmation of the conjecture proposed in [garcia2025high], namely that the nested hierarchical covariance model adequately characterizes the complex dynamics of financial instruments—here extended to the cryptocurrency market. Moreover, the empirical evidence supports a natural generalization of this conjecture to a broader class of covariance structures governed by power-law behavior.
Focusing on the specific dataset analyzed, we find that the best financial performance metrics are generally obtained using the two-step CNN-based estimator $\boldsymbol{\Xi}^{2S(\mathrm{CNN})}$ within the MVP+ portfolio framework. Notably, this result remains robust under the deliberately challenging experimental design, in which the training sample corresponds to a bull-market regime, while the out-of-sample evaluation spans a heterogeneous period beginning with a pronounced bear market. Empirical evidence further indicates that cryptocurrency returns are dominated by a strong market-wide factor, which is preserved by all covariance estimators and leads to persistently high out-of-sample volatility. As a consequence, covariance estimation primarily affects the orientation of risk—through modifications of eigenvectors and relative factor loadings—rather than the overall volatility level. Within this context, CNN-based estimators, and in particular the hybrid $\boldsymbol{\Xi}^{2S(\mathrm{CNN})}$, consistently deliver higher cumulative returns, superior Sharpe ratios, and reduced drawdowns, despite exhibiting volatility levels comparable to those of competing approaches. By contrast, the hybrid estimator $\boldsymbol{\Xi}^{\mathrm{H}}$ largely mirrors the behavior of Bitcoin itself, reflecting its strong alignment with the dominant market mode. While RMT-based and ALCA estimators improve portfolio stability or reduce turnover, they fail to match the return and drawdown performance achieved by the CNN-based methods.
@@
Nevertheless, although these results are promising in terms of adequately characterizing the complex dynamics of cryptocurrencies and in proposing a hybrid estimator that combines state-of-the-art neural network architectures with the analytical foundations of random matrix theory, they are not sufficient to achieve the maximum cumulative return. When compared with the performance of individual assets, certain instruments are capable of nearly tripling the initial capital while exhibiting a more favorable balance between risk and return, as measured by the Sharpe ratio. This indicates that, from a practical perspective, further improvements remain possible. In particular, future performance gains are likely to require the continued integration of complementary methodologies and the inclusion of an ingredient that remains relatively uncommon within econophysics—namely, asset selection and valuation grounded in fundamental analysis [graham2003intelligent]. This challenge is especially pronounced in cryptocurrency markets, where traditional fundamental metrics are largely absent.
From a statistical-physics perspective, the proposed hybrid framework extends well beyond financial applications and naturally fits within the broader class of inference problems involving low-rank deformations of Wishart ensembles. In such systems, information is not solely contained in spectral outliers but also in the collective organization and alignment of eigenvectors, which undergo sharp detectability thresholds governed by the Baik–Ben Arous–Péché (BBP) phase transition [baik2005phase, bloemendal2013limits]. Below the BBP threshold, informative factors are statistically indistinguishable from noise at the level of eigenvalues, while partial but nontrivial information persists in the eigenvector overlaps. By explicitly exploiting this eigenvector structure, the hybrid estimator effectively probes the vicinity of the BBP transition, enabling the detection of multiscale and hierarchical organization that remains inaccessible to purely eigenvalue-based denoising methods.
Competing Interests: The authors have no competing interests to declare that are relevant to the content of this article.
Data availability statement: The data that support the findings of this study are available from the corresponding author upon request.
@@
Declaration of generative AI and AI-assisted technologies in the manuscript preparation process. During the preparation of this work, the author used ChatGPT to assist with readability improvements and proofreading. After using this tool/service, the author reviewed and edited the content as needed and takes full responsibility for the content of the published article.
Author Contribution Statement. The author A.G.M. confirms being the sole contributor to this work and has approved it for publication. The author was responsible for the conception and design of the study, data collection, analysis, interpretation of results, and the writing and revision of the manuscript.